Sylvia recently had her appendix removed. Her total bill for the surgery, which was her only...

  1. Sylvia recently had her appendix removed. Her total bill for the surgery, which was her only healthcare expense for the year, came to $11,500. Her health insurance plan has a $500 annual deductible and an 80/20 coinsurance provision. The cap on Sylvia’s coinsurance share is $1,500.

How much of the bill will Sylvia pay?

How much of the bill will be paid by Sylvia’s insurance?
Answer #1

Bill Sylvia will pay = Her share of coinsurance subject to a maximum of $ 1,500 + annual deductible = min (20% x 11500, 1500) + 500 = 1,500 + 500 = $ 2,000

Bill paid by Sylvia's insurance = Total bill - payment by Sylvia = 11,500 - 2,000 = $ 9,500